# Towards a more realistic evaluation of machine learning models for bearing fault diagnosis

### v2

@@ 5 Experimental results
In this section, we apply the methodology described in Section 4 to each dataset. First, the best model architecture is identified through a CVM experiment on the UORED-VAFCLS dataset, comparing performances in the time and frequency domains. Next, we conduct evaluation experiments on all datasets, comparing three input representations across Deep Learning with two Shallow Learning models. Additionally, on the UORED-VAFCLS and PU datasets, we test different train–test split proportions to assess the impact of the number of bearings on a model’s generalization capacity.
@@ 5.1 UORED-VAFCLS dataset
The UORED-VAFCLS dataset served as the primary testbed to conduct a rigorous comparative analysis to identify the most effective deep learning architecture and input representation for bearing fault diagnosis. A comparative evaluation of five distinct 1D convolutional neural network architectures was performed using the CVM protocol on the 5 designated tuning splits. The objective was to determine the optimal combination of model, input representation (time-domain vs. frequency-domain), and hyperparameters. The results of this model selection phase, summarized in Table 7, reveal the leading performers.
Table 7: Best tuning results on the UORED-VAFCLS dataset across five architectures. Performance is reported as the mean and standard deviation of the Macro AUROC over the 5 designated tuning splits, along with the optimal hyperparameters identified.
@@
1D-ConvNet Time 85.99% +- 7.80% 64 0.001 none
Frequency 81.25% +- 6.23% 256 0.0001 none
RESNET1D Time 76.05% +- 4.06% 32 0.01 none
Frequency 87.6% +- 10.13% 32 0.001 none
CDCN Time 82.45% +- 10.54% 16 0.001 none
Frequency 89.36% +- 10.12% 32 0.0001 none
WDTCNN Time 91.09% +- 5.99% 32 0.01 none
Frequency 87.32% +- 5.12% 256 0.01 none
WDCNN Time 91.47% +- 4.87% 128 0.001 none
Frequency 93.24% +- 5.93% 256 0.0001 none
@@ 5.2 PU dataset
The experiments conducted on the curated Paderborn University (PU) dataset revealed a significant generalization challenge, with model performance approaching that of a random classifier in both the time and frequency domains. Both input representations exhibited rapid overfitting, which necessitated adding a Dropout layer with a 90% dropout rate before the output layer to regularize the model.
The final evaluation results, presented in Table 9 quantify the extent of the generalization challenge. The low mean scores and high standard deviations underscore the model’s inability to consistently learn a robust fault signature from the limited training data.
@@
WDCNN (with Dropout) Time 59.51% ± 14.14%
Frequency 61.89% ± 13.65%
Envelope Spectrum 80.16% ± 14.48%
Random Forest Time + Frequency 69.67% ± 15.73%
SVM Time + Frequency 64.42% ± 15.83%
@@
To empirically validate this diversity-limitation hypothesis, we replicated the split-proportion analysis performed on the UORED-VAFCLS dataset. As depicted in Figure 11, a clear and monotonic increase in the mean Macro AUROC was observed as the proportion of bearings in the training set increased. Although the overall performance remains low, this direct correlation provides strong evidence that the lack of component diversity is the principal bottleneck limiting model generalization on this dataset.
@@ 5.3 CWRU dataset
Envelope Spectrum 74.51 ± 9.43%
Random Forest Time + Frequency 84.40% ± 9.31%
SVM Time + Frequency 76.42% ± 9.87%
@@
Finally, we compare two distinct train-test splits on CWRU (1:2 and 2:1), following the same procedure as before, seeking to observe the impact of different split ratios on model performance. For this experiment, in order to guarantee an equal number of samples in both splits, we applied segmentation with a $12000$ size and an overlap of $53\%$, totalizing $20$ segments per signal in the 1:2 split. Figure 12 shows the results of this experiment which, although limited due to available number of configurations, is in line with those obtained for PU and UORED-VAFCLS, suggesting that increased diversity can be beneficial to performance.
@@ 6 Experiments with data leakage
Our experimental design for all datasets contrasts a leakage-free evaluation protocol with common but flawed partitioning methods that introduces data leakage.
@@ 6.1 UORED-VAFCLS dataset
  • Segmentation-level leakage: This scenario examines leakage within individual signals. While the training set only included the first 80% of each signal from the training bearings, the test set in this condition was constructed exclusively from the remaining 20% of those same signals—thus exposing the model to future portions of time series it had already partially seen during training.
@@
99.15% ± 7.19%
90.11% ± 5.42%
The results, summarized in Table 11, demonstrate the substantial and misleading performance inflation caused by both forms of data leakage. Introducing the bearing-level leakage resulted in an artificial performance increase of approximately 2.4% for time-domain inputs and 6.4% for frequency-domain inputs. The more severe segmentation-level leakage produced an higher performance increase, elevating the Macro AUROC by 7.3% and 15.3% for time and frequency representations, respectively. These findings provide compelling empirical evidence for the necessity of strict, bearing-wise data partitioning to ensure the validity and reliability of model evaluation in bearing fault diagnosis.
@@ 6.2 PU dataset
For the experimental evaluation on the PU dataset, we also generated 10 distinct data splits. In each split, the training set was constructed utilizing data from three of the four available operating conditions. From these selected conditions, 15 measurement repetitions were allocated for model training, while the residual 5 repetitions, along with all data from the fourth operating condition, were reserved to be used in the leakage scenarios. As a key modification, we truncated all training signals by removing the final 25% of each recording. Using this setup, we proceeded to test our model against three types of data leakage. First, to induce segmentation-level leakage, we followed a procedure similar to that used for the UORED-VAFCLS dataset, but now placing the 25% segments taken out of the training set into the test set (instead of 20%). The remaining types of data leakage are detailed below:
@@
98.2% ± 3.2%
100.0% ± 0.0%
The results in Table 12 reveal a similar pattern with the UORED-VAFCLS experiments, showing the impact of bearing-level leakage (with different partitioning methods) and segmentation-level leakage. Particularly, we observe the key difference between leakage across condition- and repetition-wise splits. The provided results shows that repetitions of the same experiment provide an easy way to memorize signal characteristics (a gain of 46.6% on frequency domain), sharing similar results with segmentation-level leakage, that reached perfect performance. Meanwhile, the condition-wise split also shows a high performance gain of 39.1% compared to the valid experiment, indicating that this type of data splitting also provides an easy way for model memorization.
@@ 6.3 CWRU dataset
The experimental protocol for the CWRU dataset consists in training with samples grouped by condition and fault size, as represented in Figure 13. In total, there are 12 groups, where every experiment was performed training with a single group and testing on the others. In our leakage-free protocol, the models were evaluated in groups with a different fault size, which corresponds to different bearings (e.g A $\rightarrow$ E). With bearing-level leakage, the test set consisted of the three remaining groups on each line (e.g A $\rightarrow$ B, A $\rightarrow$ C and A $\rightarrow$ D). Finally, with segmentation-level leakage, we repeated the same process as in UORED-VAFCLS, where all models were trained with only 80% of the group signals, leaving 20% separated to be used in the test set.
@@
The results are presented in Table 13 and follow the same pattern as the UORED-VAFCLS and PU datasets. In this case, both scenarios of bearing-level leakage (with condition-wise splits) and segmentation-level leakage resulted in perfect or near-perfect scores with both time and frequency domain inputs. This reinforces our hypothesis that condition-wise splits induces a strong data leakage, very similar to segmentation-level leakage on the CWRU dataset.
@@
99.8% ± 0.2%
99.9% ± 0.1%
7 Discussion and conclusion
@@
Our experimental results demonstrate that when using a leakage-free methodology, deep learning models do not exclusively yield the best performance. In fact, shallow models trained with handcrafted features can be highly competitive and, in some scenarios, even superior. This is clearly illustrated by our findings on the CWRU dataset, where a Random Forest model achieved a Macro AUROC of 84.40% ± 9.31%, outperforming all deep learning baselines. Even when not the top performer, such traditional techniques provide a crucial performance baseline. Conversely, on the PU dataset, the WDCNN with spectrum envelope input yielded the best performance (80.16% ± 14.48%), highlighting the importance of choosing a good input representation. Finally, on the UORED-VAFCLS dataset, WDCNN with frequency-domain input outperformed all other configurations, reaching a Macro AUROC of 93.12% ± 4.26%.
These findings suggest that model performance is highly dependent on both the characteristics of the dataset and the choice of representation and architecture. For example, PU showed lower and more variable results, likely due to limited variability in bearing types. In contrast, UORED’s higher scores suggest it may be a better benchmark for assessing model generalization.
@@
  • Datasets with different properties: We encourage the community to systematically test models on datasets with varying properties (e.g., PU, UORED-VAFCLS) to better assess robustness, as well as to explore multiple diversity configurations within each dataset. It is also crucial to select datasets with a structure that supports proper bearing-wise splitting and contain a sufficient number of bearings per class to enable meaningful evaluation.

### v3

@@ 5 Experimental results
In this section, we apply the methodology described in Section 4 to each dataset. First, the best model architecture is identified through a CVM experiment on the UORED-VAFCLS dataset, comparing performances in the time and frequency domains. Next, we conduct evaluation experiments on all datasets, comparing three input representations across Deep Learning with two Shallow Learning models. Additionally, we test different train–test split proportions in all datasets to assess the impact of the number of bearings on a model’s generalization capacity. We further detail the hyperparameter optimization procedure for shallow learning models, as well as the resulting best configurations for all datasets, in Appendix B. All experiments were conducted using Python with the PyTorch Lightning framework on a machine equipped with an RTX 3090 GPU and 64GB RAM.
@@ 5.1 UORED-VAFCLS dataset
The UORED-VAFCLS dataset served as the primary testbed to conduct a rigorous comparative analysis to identify the most effective deep learning architecture and input representation for bearing fault diagnosis. A comparative evaluation of five distinct 1D convolutional neural network architectures was performed using the CVM protocol on the 5 designated tuning splits. The objective was to identify the optimal combination of architecture and hyperparameters across different input representations, such as the time and frequency domains. The envelope spectrum was not included in this comparison due to computational budget constraints. The results of this architecture selection phase, summarized in Table 7, reveal the leading performers.
Table 7: Best tuning (validation) results on the UORED-VAFCLS dataset across five architectures. Performance is reported as the mean and standard deviation of the Macro AUROC over the 5 designated tuning (validation) splits, along with the optimal hyperparameters identified.
@@
1D-ConvNet Time 85.99% ± 7.80% 64 0.001 none
Frequency 81.25% ± 6.23% 256 0.0001 none
RESNET1D Time 76.05% ± 4.06% 32 0.01 none
Frequency 87.6% ± 10.13% 32 0.001 none
CDCN Time 82.45% ± 10.54% 16 0.001 none
Frequency 89.36% ± 10.12% 32 0.0001 none
WDTCNN Time 91.09% ± 5.99% 32 0.01 none
Frequency 87.32% ± 5.12% 256 0.01 none
WDCNN Time 91.47% ± 4.87% 128 0.0001 none
Frequency 93.24% ± 5.93% 256 0.001 none
@@ 5.2 PU dataset
The experiments conducted on the curated Paderborn University (PU) dataset revealed a significant generalization challenge for both time and frequency domains, whereas improved performance was observed when using the envelope spectrum. The final evaluation results, presented in Table 9, quantify the extent of the generalization challenge. The low mean scores and high standard deviations underscore the model’s inability to consistently learn a robust fault signature from the limited training data.
@@
WDCNN Time 55.08% ± 19.53%
Frequency 63.93% ± 16.06%
Envelope Spectrum 79.56% ± 13.07%
Random Forest Time + Frequency 69.76% ± 15.72%
SVM Time + Frequency 64.41% ± 15.83%
@@
Figure 11 exhibits the ROC curves obtained for inner and outer race classifiers, where we achieve an average FPR of 24.00% and 60.15%, respectively, when fixing $\text{TPR}=90\%$. From these results, we can observe that the model obtains good performance when classifying inner race faults, reaching a mean AUROC of approximately 90%. On the other hand, the outer classifier reaches a lower result of 69% with higher variance.
Figure 11: ROC curves on the PU dataset for the WDCNN trained with envelope spectrum inputs. The solid curves represent the horizontal average across 100 runs, while, for each curve, the shaded region represents one standard deviation.
@@
To empirically validate this diversity-limitation hypothesis, we replicated the split-proportion analysis performed on the UORED-VAFCLS dataset. Because the number of bearings differs across fault classes (e.g., six healthy bearings and five outer-race bearings) in this dataset, we adapted the experimental protocol to ensure exact train–test proportions. Specifically, for each split configuration, the test set was first constructed by randomly selecting the required number of bearings per health state (e.g., two bearings per class for a 3:2 ratio) and the training set was then formed by randomly sampling the corresponding number of bearings per class from the remaining pool.
As depicted in Figure 12, a clear and monotonic increase in the mean Macro AUROC was observed as the proportion of bearings in the training set increased. Although the overall performance remains relatively low, this direct correlation provides strong evidence that the lack of component diversity is the principal bottleneck limiting model generalization on this dataset.
@@ 5.3 CWRU dataset
Envelope Spectrum 74.51% ± 9.43%
Random Forest Time + Frequency 84.25% ± 8.68%
SVM Time + Frequency 75.49% ± 12.89%
@@
The ROC Curves calculated for the CWRU dataset are presented in Figure 13. We observe a similar pattern to the UORED-VAFCLS and PU datasets, where the inner classifier outperforms the others. In this case, we reach FPR = 26.39, 40.26%, and 43.05% for the inner race, outer race and ball classifiers, respectively, when fixing $\text{TPR}=90\%.$
Figure 13: ROC curves on the CWRU dataset for the Random Forest trained with handcrafted features. The solid curves represent the horizontal average across 100 runs, while, for each curve, the shaded region represents one standard deviation.
Finally, we compare two distinct train-test splits on CWRU (1:2 and 2:1), following the same procedure as before, seeking to observe the impact of different split ratios on model performance. For this experiment, in order to guarantee an equal number of samples in both splits, we applied segmentation with a $12000$ size and an overlap of $53\%$, yielding $20$ segments per signal in the 1:2 split. Figure 14 shows the results of this experiment which, although limited due to available number of configurations, is in line with those obtained for PU and UORED-VAFCLS, suggesting that increased diversity can be beneficial to performance.
@@ 6 Experiments with data leakage
Our experimental design for all datasets contrasts a leakage-free evaluation protocol with common but flawed partitioning methods that introduces data leakage. To simulate a baseline scenario in which no hyperparameters are optimized (e.g., learning rate, batch size, or the use of data augmentation), we adopted a fixed training setup for all datasets and split configurations, using a learning rate of $10^{-4}$, a batch size of 16, no normalization, overlapping training segments, and a fixed training length of 30 epochs. Because in our designs each dataset contains a very different number of training signals—approximately 24 for UORED-VAFCLS, 700 for PU, and only 3 for CWRU—we set overlap ratios of 75%, 0%, and 95%, respectively, to guarantee a sufficient number of training steps. Overall, this design aims to remove the influence of model selection as much as possible, so that any observed performance gains can be attributed exclusively to the proposed evaluation protocol, even under a fixed and non-optimized training configuration.
@@ 6.1 UORED-VAFCLS dataset
  • Segmentation-level leakage: This scenario examines leakage within individual signals. While the training set only included the first 80% of each signal from the training bearings, the test set in this condition was constructed exclusively from the remaining 20% of those same signals—thus exposing the model to future portions of time series it had already partially seen during training. [^8] The most severe scenario occurs when signal segments are randomly shuffled between training and test sets, allowing the model to observe samples from the beginning, middle, and end of the same signal during evaluation, which is also the most common form of segmentation-level leakage in the literature. Experiments that evaluate this exact configuration are available in papers such as [6, 7, 8, 9], which obtain perfect or near-perfect results.
@@
Envelope
99.94% ± 0.13%
100.00% ± 0.00%
100.00% ± 0.00%
The results reported in Table 11 correspond to the mean Macro AUROC across the 10 train-test splits for each split configuration and highlight the substantial and misleading performance inflation caused by both forms of data leakage. Introducing the bearing-level leakage resulted in an artificial performance increase of approximately 3.4%, 5.1% and 7.8% for time-domain, frequency-domain, and envelope spectrum representations, respectively. The more severe segmentation-level leakage produced a higher performance increase, elevating the Macro AUROC by 13.6%, 10.6% and 14.2% across the same representations. Notably, we observed that the performance gap between representations progressively decreased as bearing- and segmentation-level leakage were introduced, specially on the latter case, where their performance was almost equal. This result shows that invalid partitioning can distort not only performance estimates but also conclusions about the most suitable input representation during model development. Importantly, strong results were obtained even without tuning any hyperparameters, simply by introducing data leakage. These findings provide compelling empirical evidence for the necessity of strict, bearing-wise data partitioning to ensure the validity and reliability of model evaluation in bearing fault diagnosis.
@@ 6.2 PU dataset
For the experimental evaluation on the PU dataset, we also generated 10 distinct data splits. In each split, the training set was constructed utilizing data from three of the four available operating conditions. From these selected conditions, 15 measurement repetitions were randomly allocated for model training, while the residual 5 repetitions, along with all data from the fourth operating condition, were reserved to be used in the leakage scenarios. As a key modification, we truncated all training signals by removing the final 25% of each recording. Using this setup, we proceeded to test our model against three types of data leakage. First, to induce segmentation-level leakage, we followed a procedure similar to that used for the UORED-VAFCLS dataset, but now placing the 25% segments taken out of the training set into the test set (instead of 20%). The remaining types of data leakage are detailed below:
@@
Envelope
99.65% ± 0.69%
100.00% ± 0.00%
100.00% ± 0.00%
The results in Table 12 reveal a pattern similar to that of the UORED-VAFCLS experiments, showing the impact of bearing-level leakage (with different partitioning methods) and segmentation-level leakage. Particularly, we observe a significant difference between leakage across condition- and repetition-wise splits. The results show that repetitions of the same experiment provide an easy way to memorize signal characteristics (a gain of 45.5% on time domain), achieving perfect performance with frequency and envelope spectrum inputs and near-perfect performance in the time domain, indicating that repetition-wise splitting is as detrimental as segmentation-level leakage in terms of performance inflation. Meanwhile, the condition-wise split also shows a high performance gain of 30.6% (on time domain) compared to the valid experiment, indicating that this type of data splitting also provides an easy way for model memorization. Consistent with the UORED-VAFCLS protocol, the reported results correspond to the mean Macro AUROC aggregated over all train–test splits in each configuration.
@@ 6.3 CWRU dataset
The experimental protocol for the CWRU dataset consists in training with samples grouped by condition and fault size, as represented in Figure 15. In total, there are 12 groups (denoted by A–L), where each group corresponds to signals from faulty bearings located in the Drive End with the same load and fault size, yielding 3 signals per group. Note that no signals from healthy bearings are included, due to the difficulty of avoiding data leakage when they are used. With this setup, every experiment was performed by training with a single group and testing on some of the others. In our leakage-free protocol, each model was evaluated in all of the groups with a different fault size, e.g., a model trained on group A was tested on groups E, F, G, H, I, J, K, and L. With bearing-level leakage, the test set for each group consisted of the three remaining groups with the same fault size, e.g., a model trained on group A was tested on groups B, C and D. Finally, under segmentation-level leakage, we followed the same procedure adopted for the UORED-VAFCLS dataset, in which, for each signal, the first 80% of its duration was used for training and the remaining 20% was reserved for testing (note that, in this case, a model is trained and evaluated on the same group).
@@
The results are presented in Table 13 and follow the same pattern as the UORED-VAFCLS and PU datasets. In order to summarize the performance across all of the groups, the mean Macro AUROC was calculated over the test results in each split configuration. In this case, both scenarios of bearing-level leakage (with condition-wise splits) and segmentation-level leakage resulted in perfect or near-perfect scores with time, frequency and envelope inputs. Notably, a model trained using only a single signal from a single bearing for each fault type is able to achieve perfect performance when evaluated on other signals from those same bearings under different operating conditions. This reinforces our hypothesis that condition-wise splits induces a strong data leakage, very similar to segmentation-level leakage on the CWRU dataset.
@@
Envelope
100.00% ± 0.00%
100.00% ± 0.00%
100.00% ± 0.00%
While these results are clear indicators of data leakage, we now provide a deeper analysis on model failure, specifically showing how feature distributions on models without diversity are similar for signals from the same bearings. Our analysis uses the model trained on group A with time-domain inputs. This model achieves perfect AUROC for each classifier in the bearing-level leakage scenario, but shows more realistic performance with the valid (no-leakage) split: 94.8%, 53.9%, and 78.7% for inner race, outer race, and ball faults, respectively. We selected a baseline decision threshold of $0.5$ for all classifiers. Next, we extracted the model weights for each classifier and kept the top 15 values, which correspond to the features that most strongly drive failure predictions. We then selected a random 1-second segment from each signal and extracted the corresponding features. Using this weight ranking, we organized the feature distributions for all bearings and visualized the classifier output for each signal together with its group and fault type.
Figure 16: Feature-weight activation map for the model trained on group A with time domain inputs on the CWRU dataset. Rows labeled with W correspond to classifier weights, while rows labeled with F correspond to input features. The x-axis represents signals labeled by Group_FaultType_Prediction combinations, where each column shows the model response for a specific group (A-L), fault mode (IR, OR, or B), and predicted output in the same order (e.g., “110” denotes that the input was diagnosed as containing IR and OR faults).
Figure 16 shows our results in detail and highlights several important points. First, we clearly observe that a model trained with a given bearing produces similar feature distributions for other signals from that same bearing. In this experiment, groups that share the same bearing between training and testing but differ in operating conditions—such as B, C, and D—form blocks of similar features. These feature blocks are also associated with the highest weights of each classifier, indicating that the model learned to prioritize them when detecting faults. However, other bearings (such as the inner and outer race in groups E to H) show inconsistent values for these same features, suggesting that they are not true fault indicators but rather spurious correlations with the bearings seen during training. For ball faults, feature distributions appear more consistent across bearings, since most ball fault signals activate similar high-importance features. Even so, these same features are also triggered by bearings with other fault types in groups E to H. Overall, these results suggest that a model trained on only a few bearings tends to memorize intrinsic signal characteristics instead of learning fault-related indicators. Evaluating such models on signals from the same bearings is therefore close to measuring training performance rather than true generalization.
7 Deployment considerations
Considering the importance of a realistic, leakage-free evaluation highlighted by the previous experiments, we encourage researchers to apply our methodology to their own datasets. In many industrial scenarios, implementing a bearing-wise split may not be straightforward, especially when bearing identifiers are not explicitly available and data acquisition is typically organized by sensor location rather than by individual bearings. A common setup involves multiple sensors mounted at different points on a rotating machine (e.g., a conveyor belt), with each sensor positioned close to a bearing. In such cases, a sensor-based split, in which models are evaluated on data from unseen sensors, would be comparable to a bearing-wise split.
Although this approach adapts the bearing-wise split to realistic industrial settings, it may still be susceptible to data leakage, particularly because sensors can capture vibrations originating from nearby bearings (e.g., bearings located in side positions). When possible, we therefore recommend more robust strategies to promote proper generalization, such as splitting data by groups of nearby sensors (which are attached to individual bearings or other components) or by entire machines when multiple machines are available. This requirement extends naturally to online learning scenarios, where model updates must be performed without using data from the evaluation group in order to avoid introducing data leakage.
With these progressively stricter data partitioning strategies, measured model performance can more realistically reflect deployment conditions, e.g., if a model is intended for operation on unseen machines, a machine-wise split would be the appropriate choice. Nevertheless, machine-wise splitting can be challenging due to domain shift, which may require the use of domain adaptation or domain generalization techniques. A similar issue arises when models are trained on public benchmark datasets and subsequently tested on data from different machines, a scenario that likewise characterizes domain shift comparable to machine-wise partitioning.
Finally, although this work focuses on bearing fault diagnosis, our multilabel framework can be extended to include multiple fault modes (e.g., lack of lubrication), allowing a specific operating point to be defined for each classifier and enabling an evaluation methodology that is robust to class prevalence, which is particularly useful in condition monitoring applications.
8 Conclusions
@@
Our experimental results show that, under a leakage-free methodology, the deep learning models we were able to evaluate within our computational budget and development time did not consistently achieve the best performance. In several cases, shallow models trained with handcrafted features produced highly competitive results. This is evident in the CWRU dataset, where a Random Forest model achieved a Macro AUROC of 84.25% ± 8.68%, outperforming all deep learning baselines. Even when they are not the top-performing approach, such traditional techniques provide an important performance reference. Conversely, on the PU dataset, the WDCNN with spectrum envelope input yielded the best performance (79.56% ± 13.07%), highlighting the importance of choosing a good input representation for the dataset at hand. Finally, on the UORED-VAFCLS dataset, WDCNN with frequency-domain input outperformed all other configurations, reaching a Macro AUROC of 93.12% ± 4.26%.
Another conclusion that can be derived from our results is the importance of ensuring sufficient bearing diversity in the training set, as increasing the number of distinct bearings consistently led to improved performance across all evaluated datasets. Importantly, our experimental protocol demonstrates that simply enlarging the training set is not sufficient when this increase does not introduce new bearings, since all split configurations (1:4, 2:3, 3:2 and 4:1) were trained with the same number of samples (the total number of samples seen by the model during the entire training is kept fixed) and yet exhibited substantially different performance depending on the number of bearings available for training. This behavior indicates that bearing diversity, rather than sample count alone, plays a critical role in achieving generalizable performance. Taken together, these findings suggest that model performance is highly dependent on both the characteristics of the dataset (e.g. sufficient bearing diversity) and the choice of model family and input representation.
@@
  • Datasets with different properties: We encourage the community to systematically test models on datasets with varying properties (e.g., PU, UORED-VAFCLS) to better assess robustness, as well as to explore multiple diversity configurations within each dataset. It is also crucial to select datasets with a structure that supports proper bearing-wise splitting and contain a sufficient number of bearings per class to enable meaningful evaluation (which is not the case of the CWRU).